A dealer wishes to obtain 80 pounds of mixed cookies to sell for one dollar per pound. If he mixes cookies worth $1.20 per pound with the cookies worth $.70 per pound, find the number of pounds of each kind he should use
Answer by fractalier(6550)

Call the amount at $1.20, x.
Then the amount at $.70 would be 80-x.
The set up is then
1.20x + .70(80-x) = 1(80)
1.2x + 56 - .7x = 80
.5x = 24
x = 48 pounds of the $1.20
80-x = 32 pounds of the $.70